A, B and C are on a betting game. B loses Php 350 of his money to A. As a result, A now has twice as much as what is left with B. Then, C loses Php 700 to B. As a consequence, C now has only one-third as much money as B would then have. If A has loses Php 210 to C, C will have as much as money as A would have left. How much did each have at the start?


1
Expert's answer
2022-04-26T01:01:14-0400

"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 A & B & C \\ \hline
 X & Y & Z \\
 \hdashline
 X+350 & Y-350 & Z
\\ \hdashline
X+350&Y+350&Z-700
\\ \hdashline 
X+140& Y+350&Z-490
\end{array}"


"B" loses Php 350 of his money to "A" : "A" has "X+350" and "B" has "Y-350."

"A" now has twice as much as what is left with "B": "X+350=2(Y-350)."


"C" loses Php 700 to "B" : "C" has "Z-700" and "B" has "Y+350" .

"C" now has only one-third as much money as "B" would then have: "Z-700=\frac{1}{3}(Y+350)."


If "A" has loses Php 210 to "C" , "C" will have as much as money as "A" would have left: "A" has "X+140" and "C" has "Z-490" and "X+140=Z-490."


"\begin{cases}
X+350=2(Y-350)
\\
Z-700=\frac{1}{3}(Y+350)
\\
X+140=Z-490
\end{cases}" "\begin{cases}
X=2Y-1050
\\
Y=3Z-2450
\\
Z=X+630
\end{cases}" "\begin{cases}
X=6X-2170
\\

Y=3X-560
\\
Z=X+630
\end{cases}" "\begin{cases}
X=434
\\ Y=742\\ Z=1064\end{cases}"


Answer: "A" - "434," "B" - "742," "C" - "1064."